OpenStudy (anonymous):

A fair four-sided die with its four faces number 1 to 4 are tossed together with a fair coin. What is the chance that an even number and a head appear together?

OpenStudy (anonymous):

even number is {2,4} out of 4 possibilities. so probability of even number is 1/2. evidently if you toss a coin the probability of heads is also 1/2 now multiply to get your answer \[\frac{1}{2}\times \frac{1}{2}=\frac{1}{4}\]